OpenStudy (anonymous):

need help with system: -14=-20y-7x 2x=10y +4 I have 20y-4x=-8 and -20y-7x=-14, but i think i did something wrong after that

OpenStudy (anonymous):

what did you do wrong?

OpenStudy (anonymous):

i got -11x=-24

OpenStudy (anonymous):

but I got x = 2 and y = 0 :-(

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

-20y + 4x = 8 => =20y = 8 -4x - 20y - 7x = -14 -8 - 4x - 7x = -14 -11x = -22 x = 2 -20y + 4(2) = 8 -20y = 0 y = 0

OpenStudy (anonymous):

oh, i see